OpenStudy (anonymous):

Did I get it correct? The area of a rectangle is the product of width and length. If a rectangle has length 18m√5m and width √4m , which of the following is the area of the rectangle? A.36m^2√5 B.72m^2√5 C.72m√5m <-----My Answer D.36m√5m

OpenStudy (earthcitizen):

uhm, A

OpenStudy (anonymous):

The square root of 4 is 2, so the product of length and width is\[18\sqrt{5}m*2m = 36m^{2}\sqrt{5}\]

OpenStudy (anonymous):

Ah I read the wrong thing in m notebook I got to 18√20 after 18√5m√2m and ended up doing 18√10*2 then 18*2 = 36, then it was 36√10 or 36√5*2, 36*2 = 72√5m lol but thanks

hero (hero):

@EarthCitizen, just because the OP offered an answer, that doesn't mean give the answer. You still have to provide an explanation

OpenStudy (earthcitizen):

no pp, @Hero

OpenStudy (anonymous):

yea u did!! good job